# Personality Structured Interview for Large Language Model Simulation in Personality Research

### v1

@@ 6 Conclusion
In summary, this study proposes a novel method for LLM to simulate human personality distributions—PSI and provides a comprehensive explanation of its development process and corresponding dataset.
Through three experiments, we validated PSI’s effectiveness. Results show that PSI matches or even outperforms two-hour interviews (Park et al., 2024) in personality simulation and surpasses existing methods in terms of both item-level accuracy and overall psychometric indicators. One experiment also examined PSI’s ability to simulate personality-related behavior, revealing that although LLMs approach human-like performance, they still exhibit idealized tendencies.
In summary, a theory-informed structured interview like PSI can better simulate human-like psychometric data. We explore the role of theory-informed structured interviews in simulation in greater detail and examine their potential to advance research, both of which are elaborated upon in Appendix C.
@@
This study has several limitations. First, we explored the potential of PSI in advancing research. Our single experiment demonstrates that theory-informed questions can effectively elicit information embedded in human narrative responses—–particularly those relevant to the target personality construct—–allowing LLMs to simulate the corresponding behavior. However, additional experiments are needed to investigate its broader applications.
Second, the personality assessment in this paper relies on self-reported personality scales, which can be viewed as a limitation of the study. However, these self-report scales are specifically designed for humans to target personality traits across individuals, in ways that cannot be directly located within LLMs themselves (see Appendix C for a detailed discussion on psychometrics). Furthermore, self-report scales are used often because respondents have privileged access to their own personality, and even with their well-known limitations (e.g., faking or social desirability in self-reports; see Appendix I.3 for the social desirability analysis), they remain valuable to evaluate the personality traits exhibited by LLMs. However, because the primary goal of this study is to have the LLM simulate human responses and behaviors that correspond to extracted personality information, it is reasonable to assess LLMs using methods designed to measure human personality traits.
Moreover, this paper tests only two types of LLMs, which may somewhat limit its scope, given the array of LLMs that are now available. However, the conclusions drawn from our two models were highly consistent, and we have reason to believe that the advantages of the PSI method are generally applicable across different LLMs. Furthermore, we designed three different experiments and conducted comprehensive psychometric analyses of various methods within LLMs, ensuring greater rigor and reliability of the results.
@@
We hereby confirm that all authors of this study are aware of the provided ACL Code of Ethics and comply with the Code of Conduct.
Human Sample Data
@@
This paper discusses the comparison between multiple sets of human data and results generated by LLMs. The collection of human data strictly adhered to relevant ethical guidelines and received approval from the Institutional Review Board. To ensure fair treatment of participants and proper recognition of their contributions, reasonable compensation or course credit (for student participants) was provided.
Throughout the research process, we placed a strong emphasis on transparency and openness, also ensuring that all participants signed informed consent forms. Additionally, to safeguard data privacy and uphold ethical standards, the publicly available dataset underwent rigorous screening to exclude any personally identifiable information and was shared only with explicit public consent.
Appendix A Appendix: The Structure of Personality
A key question in personality-related LLM research pertains to personality structure: What is the nature and breadth of the personality traits we want to simulate the human personality distributions? In the research literature, personality structures often emerge from applying factor analysis to individuals’ responses to a large number of personality-relevant items. This approach is what has been used to identify the five personaltiy factors in the FFM. Moreover, personality is better understood in terms of one’s continuous standing on each of multiple dimensions rather than as static types or profiles. Research data clearly supports this view (Wilmot, 2015; Wilmot et al., 2019). Dividing individuals into limited categories (e.g., 16 types in MBTI) artificially segments continuous dimensions into discrete units, which may overlook important individual differences (Ones and Wiernik, 2018).
Another question is whether we should incorporate personality traits and structures beyond the FFM to achieve a more comprehensive understanding of personality, given that additional personality variables and alternative structures have been proposed over the years. It turns out that there can be conceptual overlap among these models (e.g., Hough et al., 2015).
For example, a meta-analysis by Joseph and Newman (2010) revealed that emotional intelligence (EI) shows statistically and practically significant relationships with neuroticism and extroversion within the FFM. In fact, when controlling for personality variables, the unique contribution of EI almost disappears. Similarly, Credé et al. (2017) conducted a meta-analysis on grit and found that its core components can largely be explained by conscientiousness, with little added predictive validity beyond that. Moreover, although the HEXACO structure introduces an honesty-humility dimension, the remaining five dimensions align closely with the FFM structure (Lee and Ashton, 2004, 2006). Research by Cutler and Condon (2023) further indicates that nearly all personality semantic information can be classified within the FFM structure.
These studies collectively suggest that although further subdivision of personality structures might provide new perspectives, it often leads to conceptual redundancy and measurement complexity without necessarily enhancing predictive validity or theoretical value. Rather than pursuing a wide range of personality frameworks, we will delve deeper into the FFM structure, which is very widely accepted by personality psychologists.
Appendix B Appendix: Personality Structured Interview Questions
The initial draft of the question pool was created and refined through collaborative discussions among subject matter experts (SMEs). The SME team consisted of two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology. To ensure the quality and clarity of the questions, pilot testing was conducted with six undergraduate research assistants from a personality research lab. This iterative process of development and feedback led to the construction of the personality structured interview.
The development process is similar to the development of psychological tests or scales (psychometric). We have provided more details on psychometrics and the development framework in Appendix C.
# Questions
2 Thinking back, what kind of student were you in school?
4 What was your favorite subject in school, and why?
5 What was your least favorite subject in school, and why?
6 Still thinking back, who were your heroes when you were young and why?
7 When you were little, what did you want to be when you grew up? And why?
9 If you had complete freedom, what would your dream job be, and why?
10 How have your dreams and goals changed throughout your life?
12 How are you similar or different from your parents and/or siblings?
14 What was the best part of your childhood?
15 What do you think were the worst parts of your childhood?
18 What do you do now for a living? And why did you choose it?
19 Please describe your typical work day.
20 What is the best and worst part of your current work?
22 Moving on, what are your adult friendships like?
27 What hobbies or other interests do you have?
28 What things frighten you now?
29 What were some things that frightened you most as a child?
Table 3: Structured Interview Questions
Appendix C Appendix: Psychometrics and Structured Interview Development Framework
Psychometrics is a field of psychology dedicated to the theory and practice of psychological measurement. It primarily focuses on quantifying psychological traits, behaviors, and abilities through systematic testing and analysis. Psychological traits, such as personality dimensions, cognitive abilities, and emotional states, are inherently abstract constructs that cannot be measured directly. Therefore, psychometricians rely on tools like surveys, questionnaires, scales, or structured interviews to infer these traits through observable indicators or responses.
Measuring Psychological Traits
These behaviors/identities are translated into measurable items (for scale; e.g., “I enjoy being the center of attention”) or questions (for interview; e.g., “What are your strongest qualities as a friend? In other words, what makes you a great friend to have?”). The challenge lies in ensuring that these items/questions accurately and consistently capture the construct across different populations and contexts.
Theory-Informed Structured Interview for LLM Data Simulation
The Potential for Advancing Research
Structured Interview Development Framework
# Steps
2 Prepare a set of structured interview specifications—structured interview blueprint.
3 Build an initial question pool.
4 Have questions reviewed by substantive experts (and revised as necessary).
5 Hold preliminary question tryouts.
9 Develop guidelines for administration, and interpretation of the structured interview.
Table 4: Structured Interview Development Framework
Prepare a set of structured interview specifications—structured interview blueprint: A blueprint outlines the structured interview’s structure and content, specifying how questions will be distributed across the construct’s dimensions or components. It typically includes information on the number of questions per domain, questions content.
Appendix D Appendix: Personality Structured Interview Dataset and Data Collection
Institutional Review Board approval was obtained. The data was collected through an online questionnaire, followed by an online structured interview.
Since we are still actively collecting data, we will share a part of the dataset that does not include any personal privacy information and has received permission for data sharing. Each example is composed of the following characteristics:
  1. 1.
    Gender: The gender of the participant.
  2. 2.
    Race: The racial background of the participant.
  3. 3.
    English: Whether English is the participant’s first language.
  4. 4.
    Age: The participant’s age.
  5. 5.
    Weight: The participant’s weight.
  6. 6.
    Height: The participant’s height.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
    Item1–Item60: Responses to each item of the BFI-2. Refer to Appendix E for item descriptions and scoring guidelines.
Appendix E Appendix: BFI-2 Scale
@@
Table 5: BFI-2 Scale
Scoring: Reverse-keyed items are denoted by “R.”
In psychological measurement and scale development, reverse coding is a common and essential technique. Its primary purpose is to ensure that the scoring direction of all items remains consistent, thereby improving the reliability of the measurement results and the accuracy of their interpretation.
A balanced arrangement of positive and negative items helps reduce response biases, such as consistency effects or response patterns where participants select the same option repeatedly. These tendencies can mask the respondent’s true attitudes or behavioral traits, leading to distorted measurement outcomes. By alternating positive and negative statements and applying reverse coding to the appropriate items, this issue can be effectively mitigated.
Reverse coding can also help reduce the influence of social desirability bias. When all items are presented in the same direction, participants may easily guess the purpose of the test and provide responses that align with perceived expectations. By mixing positive and negative statements and applying reverse coding, the scale can disrupt this pattern, making it harder for participants to determine the “correct” answers, thus providing a more genuine reflection of their inner states or attitudes.
@@
Table 6: BFI-2 Domain Level with Item Numbers
@@
Table 7: BFI-2 Facet Level with Item Numbers
Appendix F Appendix: Prompts List
Personality Scale Prompt Format
Table 8 presents the prompts used to generate LLM responses for the selected personality test on a Likert scale, where personality_description denotes the personality prompt.
In the PSI method, the personality prompt integrates both the interview question and the interviewee’s corresponding response, as collected in the PSI dataset (see Appendix D). For the Persona and Shape methods, more detailed descriptions of the personality prompts can be found in Appendix G.
The test_item corresponds to each individual item in the BFI-2 scale.
For the following task, respond in a way that matches this description: {personality_description}.
Considering the statement, please indicate the extent to which you agree or disagree on a scale from 1 to 5 (where 1 = “disagree strongly”, 2 = “disagree a little”, 3 = “neither agree nor disagree”, 4 = “agree a little”, and 5 = “agree strongly”): {test_item}.
Table 8: Prompt format for gathering LLMs’ responses to BFI-2 scale.
Behavioral Question Prompt Format
The complete prompt format for eliciting LLMs’ responses to personality-related behavioral questions is defined in Table 9.
The questions_and_responses refers to the transcript of a structured human personality interview, while question_list comprises statements evaluating OCB and CWB, as detailed in Table 10 and Table 11 in Appendix G. The model is instructed to rate each statement using a frequency scale from one to five.
Task: Simulate an individual’s behavior and predict their responses to a series of work-related questions.
Data to Analyze: {questions_and_responses}
Instructions:
1. Simulate the Individual: Use the provided data to simulate the individual’s personality, behavior, and work habits.
2. Predict Responses:
- Based on the provided data, simulate the individual and answer the question.
- Recall that in the past year, how often have you done the following things? (1 = Never, 2 = Once or twice, 3 = Once or twice per month, 4 = Once or twice per week, 5 = Every day)
Questions: {question_list}
Output Format:
Question Score: [Provide the numerical score here.]
Appendix G Appendix: Additional Experiment Settings
Pearson Correlation Coefficient
The correlation calculation formula is as follows:
$X_{i}$ and $Y_{i}$ represent the data values for LLMs and human. The symbols $\overline{X}$ and $\overline{Y}$ represent the means of variables $X$ and $Y$, respectively. The numerator, $\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})$, represents the covariance between $X$ and $Y$. The denominator, $\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}$, standardizes the result, constraining the value of $r$ to range between -1 and 1.
Persona Method
The Persona method is based on the Persona-Chat dataset constructed by Zhang et al. (2018). The dataset consists of persona descriptions, and each is made up of five short sentences containing demographic information collected through Amazon Mechanical Turk crowdsourcing. To avoid sentence similarity or repetition, these persona descriptions were required to be rewritten (e.g., changing “I am very shy” to “I am not a social person”). Zhang et al. (2018) demonstrated through machine learning model validation and human evaluations that such persona descriptions provide an effective method to enhance personalization. Currently, incorporating personal profiles into prompts is widely used in research related to LLM agents (Park et al., 2023; Wang et al., 2023; Xi et al., 2023).
In the current study, we treated each persona description as an individual entity (i.e., a single subject) and randomly selected 300 persona descriptions from the dataset. One example is, “I wear a lot of leather. I have boots I always wear. I sleep in late during the day. I listen to metal music. I have black spiky hair.”
Shape Method
The Shape method is based on the work of (Serapio-García et al., 2023), who introduced a prompting approach to shape synthetic personality in LLMs along desired dimensions. The researchers expanded upon Goldberg (1990) lexical hypothesis, expanding his list of 70 bipolar adjectives (Goldberg, 1992) to include 104 trait adjectives. Additionally, they employed linguistic qualifiers commonly used in Likert-type scales (Likert, 1932), such as “a bit,” “very,” and “extremely,” to set target levels for each adjective. This resulted in a fine-grained prompting method with nine levels: 1. extremely low adjective; 2. very low adjective; 3. low adjective; 4. a bit low adjective; 5. neither low adjective nor high adjective; 6. a bit high adjective; 7. high adjective; 8. very high adjective; 9. extremely high adjective.
In our study, each prompt involves five randomly selected adjective markers from a specific personality domain. These markers are positioned after a consistent linguistic qualifier to set the prompt at one of nine intensity levels. For example, one prompt is: “You are extremely friendly, extremely energetic, extremely assertive, extremely bold, and extremely active.” These are five positively-keyed adjectives targeting extraversion. In this case, the prompt seeks to create a highly sociable and dynamic personality profile, which might result in responses characterized by enthusiasm, confidence, and proactivity. We also randomly select 300 prompts here.
CFA Model
The basic form of the CFA model is:
$y=\Lambda\eta+\epsilon$
Model Fit Information
$\chi^{2}=(N-1)\times F_{\text{ML}}$
Degrees of Freedom, df: The df represent the relationship between model parameters and observed variables:
$df=\frac{p(p+1)}{2}-q$
$\text{RMSEA}=\frac{\chi^{2}-df}{df(N-1)}$
Based on Hu and Bentler (1999), CFI/TLI $\geq$ .95, RMSEA $\leq$ .06, and SRMR $\leq$ .08 are considered good fit thresholds.
Tucker’s Congruence Coefficient
The coefficient ranges from -1 to 1, where values close to 1 indicate high similarity (congruence) between the factor solutions, values close to 0 indicate low similarity, and negative values indicate a dissimilar or inverse relationship. According to Lorenzo-Seva and Ten Berge (2006), a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity. However, this is a relatively lenient criterion; specific differences still need to be determined based on the factor loadings.
Behavior Variable Measures
@@ Appendix G Appendix: Additional Experiment Settings
Organizational Citizenship Behavior (OCB): OCB was measured using ten items from Spector and Fox (2010) to assess extra-role behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: “In the past year, how often have you helped new employees get oriented to the job?”. Internal consistency was Cronbach’s alpha = .83.
@@
Counterproductive Work Behavior (CWB): CWB was measured using ten items from Spector and Fox (2010), designed to assess harmful workplace behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: “In the past year, how often have you ignored someone at work?”. Internal consistency was Cronbach’s alpha = .86.

### v2

@@ 5 Conclusion
This study introduces a novel method, PSI, for simulating human personality data with LLMs and provides a detailed account of its development and the associated dataset. Across three experiments, we evaluated the effectiveness of PSI. The results show that PSI performs well in simulating personality, yielding moderately strong and statistically significant correlations with human self-report data. Moreover, when modeling the distribution of personality data at the population level, PSI outperforms existing methods in both human-like heterogeneity and broader factor analysis results.
Our experiments also assessed PSI’s ability to simulate personality-related behavior. Although LLMs approximate human-like patterns, they still exhibit idealized responses that deviate from the natural variability found in human behavior. In summary, theory-informed structured interviews such as PSI offer a more realistic and psychometrically grounded approach to simulating human-like data. We further discuss the utility of such data in simulation-based research and its potential to improve future psychometric studies in Appendix G.
@@
This study is not without limitations. First, our initial evidence suggests that theory-driven prompts can successfully elicit personality-relevant information embedded in human narratives, enabling LLMs to simulate corresponding psychometric data. However, further research is necessary to examine the robustness and generalizability of this approach across diverse contexts and constructs.
Second, the personality assessment in this study is based on self-report measures originally designed for humans. These tools are intended to capture the stable, internal personality that individuals can access and report, which LLMs, by design, do not inherently possess (see Appendix G for a detailed discussion on psychometric considerations). Nevertheless, LLMs can simulate human-like responses that reflect particular personality patterns when guided by theory-informed prompts. Because the goal of this study is not to assert that LLMs have personality, but rather that they can simulate personality-driven behavior, it is reasonable to apply human-based assessment tools to evaluate the quality and fidelity of that simulation.
@@
This paper presents a comparison between human data and LLM-simulated data by three different methods. All human data were collected in strict accordance with relevant ethical guidelines and were approved by the Institutional Review Board. Participants received either reasonable monetary compensation or course credit (in the case of student participants) to ensure fair treatment and appropriate recognition of their contributions.
We placed a strong emphasis on transparency and ethical integrity throughout the research process. All participants provided informed consent prior to participation. Moreover, to protect privacy and maintain ethical standards, the publicly shared dataset was carefully screened to remove any personally identifiable information.
Acknowledgments
This research project is partially supported by the Science of Trustworthy AI Award from Schmidt Sciences, the AI2AI Award from Amazon, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, and by the National Science Foundation (NSF) under Award Number 2522411. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NSF.
Pengda Wang: conceptualization (equal, lead), methodology (equal, lead), software (lead), formal analysis (lead), investigation (lead), data curation (lead), writing - original draft (lead), writing - review & editing (equal, lead), visualization (lead); Huiqi Zou: methodology (support), validation (support), formal analysis (support), investigation (support), data curation (support), writing - original draft (support), writing - review & editing (equal, support), visualization (support); Han Jiang: methodology (support), validation (support), investigation (support), data curation (support), writing - original draft (support), writing - review & editing (equal, support), visualization (support); Hanjie Chen: writing - review & editing (equal, support); Tianjun Sun: conceptualization (equal, lead), methodology (equal, lead), validation (support), formal analysis (support), investigation (support), resources (equal, lead), data curation (support), writing - original draft (support), writing - review & editing (equal, lead), supervision (equal, lead), funding acquisition (equal, lead); Xiaoyuan Yi: review & editing (equal, support), supervision (equal, lead); Ziang Xiao: conceptualization (equal, lead), methodology (equal, lead), validation (support), formal analysis (support), investigation (support), resources (equal, lead), data curation (support), writing - original draft (support), writing - review & editing (equal, lead), supervision (equal, lead), funding acquisition (equal, lead); Frederick L. Oswald: methodology (support), writing - review & editing (equal, lead).
Appendix A Appendix: BFI-2 Scale and Behavioral Measures Scales
A.1 BFI-2 Scale
@@
Table 2: BFI-2 scale.
Scoring: Reverse-keyed items appear as “R.”
Reverse coding is a widely used technique in psychological measurement and scale development. Its primary purpose is to align the scoring direction of all items, ensuring consistency across the scale. This process enhances the reliability of the overall measurement and supports a more accurate interpretation of the results.
A balanced inclusion of positively and negatively worded items is a well-established strategy to reduce response biases—such as acquiescence bias, consistency effects, or patterned responding (e.g., selecting the same scale point across items). These biases can obscure respondents’ true attitudes or behavioral tendencies, compromising measurement validity.
Although it is not guaranteed, reverse coding can help reduce the influence of social desirability bias. When all items are presented in the same direction, participants may easily guess the purpose of the test and provide responses that align with perceived expectations. By mixing positive and negative statements and applying reverse coding, the scale can disrupt this pattern, ideally making participants reflect more carefully on their inner states or attitudes when providing responses.
@@
Table 3: BFI-2 domain level with item numbers.
@@
Table 4: BFI-2 facet level with item numbers.
A.2 Behavioral Measures Scales
@@
Organizational Citizenship Behavior (OCB): OCB was measured using ten items from Spector and Fox (2010) to assess extra-role behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: ‘‘In the past year, how often have you helped new employees get oriented to the job?’’. Internal consistency was Cronbach’s alpha = 0.83.
@@
Counterproductive Work Behavior (CWB): CWB was measured using ten items from Spector and Fox (2010), designed to assess harmful workplace behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: ‘‘In the past year, how often have you ignored someone at work?’’. Internal consistency was Cronbach’s alpha = 0.86.